# CIC-Trap4Phish: A Unified Multi-Format Dataset for Phishing and Quishing Attachment Detection

### v2

@@ I Introduction
The remainder of this paper is organized as follows. Section II reviews existing works and related datasets. Section III describes the steps involved in generating the proposed dataset, including data collection. Section 4 explains the processes of feature extraction, feature selection, and classification. Section IV presents the experimental results and evaluation criteria. Finally, Section V concludes the paper and outlines future research directions.
@@ III-E QR code
IV Experimental Results and Evaluation
@@ IV-A Evaluation Metrics
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
$\text{Precision}=\frac{TP}{TP+FP}$ (2)
$\text{Recall}=\frac{TP}{TP+FN}$ (3)
$F1\text{-score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
IV-B Evaluation Results
@@
IV-B1 Word Doc Evaluation
In the feature extraction step, we extracted 43 static features from Word Document files, as described in section 4. To obtain a compact set of efficient and discriminative features, we selected top 10 features from SHAP and Random Forest feature importance rankings. These selected features were evaluated using three classifiers, and the detailed results are presented in Table III. Additionally, Figure 8 (a) provides a visual representation of the results. The selected set of 10 features achieved 100% Accuracy, Precision, Recall, and F1-score, demonstrating the strong effectiveness of the selected feature subset. The confusion matrices for all three classifiers are shown in Figure 5. These matrices confirm the near-perfect classification performance of benign vs. malicious Word Document files by considering the top 10 selected features. Random Forest and XGBoost achieved perfect classification with no misclassifications, while the Decision Tree confusion matrix shows there is only a negligible error rate. These results confirm that the top 10 selected features enable highly accurate and efficient detection of malicious Word document files.
@@ IV-B1 Word Doc Evaluation
IV-B2 Excel Evaluation
In order to evaluate the extracted and selected Excel features, we performed the set of experiments that were conducted on Word documents. As explained before, a top 10 features were selected out of 48 extracted features and then the three machine learning classifiers were used for evaluation phase. Figure 8 (b) illustrates the four evaluation metrics, while Table III presents their detailed results. Additionally, Figure 6 shows the confusion matrices of all three classifiers as well, which demonstrates the effectiveness of the extracted features.
@@ IV-B2 Excel Evaluation
IV-B3 PDF Evaluation
For PDF documents, a total of 40 features were initially extracted, as discussed in Section 4. Subsequently, the top 10 features were selected based on their importance scores. Similar to the previous file formats, these top 10 features were evaluated using the three classifiers described earlier. Table III presents the detailed evaluation metrics. The results show that the selected features are highly effective, enabling accurate and efficient classification of benign and malicious PDF files. The results of these three models are provided in Figure 8 (c), which shows that all four evaluation metrics, Accuracy, Precision, Recall, and F1-score, exceed 99%.
@@ IV-B3 PDF Evaluation
IV-B4 HTML Evaluation
From each benign and malicious HTML sample, a set of 40 features were initially extracted, and subsequently, 13 features were selected that provided the best Accuracy, Precision, Recall and F1-score, while providing low training and testing time. By applying three classification models, XGBoost and Random Forest classifiers achieved approximately 93% in Accuracy, Precision, and F1-score, with a Recall of 92% which show the strength of selected features. The Decision Tree classifier achieved around 89% across all metrics, which still indicates satisfactory classification performance. The detailed results of these models for the selected HTML features are presented in Table III and visualized in Figure 8 (d). The confusion matrices for all three classifiers are also illustrated in Figure 9.
Table III: Per-classifier performance by data type.
@@ IV-B4 HTML Evaluation
IV-B5 QR code Evaluation
In order to evaluate the QR codes samples that were already generated from the malicious and benign URLs, we firstly followed an image processing approach by applying a basic CNN model to the QR code images. The corresponding performance results are provided in Table IV. As shown in this table, the CNN model accuracy is moderate, despite being trained on a large dataset, which proves that the image-based approach has some limitations in detection phase. To further investigate these limitations, additional visual and structural similarity analyses were performed. The results of these analyses are presented in Figure 10.
@@ IV-B5 QR code Evaluation
The next step that we followed involved analyzing the URL strings corresponding to the QR codes to evaluate the effectiveness of benign and malicious URL samples in the detection phase from a lexical perspective. In order to examine different models, several lightweight language models (LLMs), including BERT-Tiny, DeBERTa-v3, ModernBERT, and DeepSeek-R1, were applied to the URL dataset in this experiment. The results presented in Table IV proves that LLM-based methods (URL lexical analyses) achieve better performance compared to the CNN model, as reflected by their higher Precision, Recall, and F1-score values. Moreover, the training time of these models is significantly lower than image based approach, since they process character level representations and tokenized text rather than images.
@@
Table IV: Performance results for malicious QR code detection using image-based approaches and recent lightweight LLMs applied to URL-based analysis.
Detection Classifier Approach Train Time (s) Inference Time (s) Dataset Size Precision (M) Recall (M) $\boldsymbol{F_{1}}$ Score (M)

### v3

@@ I Introduction
The remainder of this paper is organized as follows. Section II reviews existing works and related datasets. Section III describes the steps involved in generating the proposed dataset, including data collection. Section 4 explains the processes of feature extraction, feature selection, and classification. Section V presents the experimental results and evaluation criteria. Finally, Section VI concludes the paper and outlines future research directions.
@@ III-E QR code
IV Feature Engineering and Classification
Our study focuses on five data types, including Word documents, Excel spreadsheets, PDF files, HTML pages, and QR codes. For the first four file formats, we developed a static feature-based approach, which includes the feature extraction and feature selection phases, followed by the use of traditional machine learning models for the detection phase. This approach enables the development of lightweight and efficient detection models, as it does not rely on dynamic file execution. In contrast, QR codes were analyzed directly using a Convolutional Neural Network (CNN), enabling image-based learning without explicit feature extraction. To further enhance detection performance, a URL lexical analysis was introduced using lightweight transformer models after decoding the URLs embedded within the QR codes.
The development of effective machine learning models for detecting malicious attachments across four data types requires a dedicated feature extraction pipeline capable of capturing both general and format-specific characteristics. The objective was to extract structural properties, embedded content, and behavioral patterns through static analysis methods that do not require file execution. A customized parser was implemented for each document type to extract diverse features, including file metadata, macro indicators, script patterns, embedded object structures, and related attributes.
After feature extraction, the set of extracted features went through a two-step feature selection process to identify the most important attributes for classification tasks and also reduce dimensionality in order to have more efficient training. The SHAP (SHapley Additive exPlanations) model was employed to quantify the contribution of each feature to the model’s prediction outcomes, providing interpretability through feature impact analysis. In parallel, Random Forest feature importance scores were computed to rank attributes based on their contribution to decision splits within the ensemble. The integration of these two evaluation methods allowed us to choose a smaller set of essential features for each file type. The selection process maintains only the most important features which leads to improved detection results and better model generalization capabilities.
IV-A Word
In order to extract high level features from Word documents, we created a static feature extraction pipeline which focuses on both high-level and structure-specific characteristics of .docx files. The detection system was designed to identify essential malicious indicators, including embedded macros, suspicious VBA keywords, Dynamic Data Exchange (DDE) links, and abnormal XML structures that appear in harmful Word document attachments.
The extraction process produces 43 features for each file. The extracted features consist of five main categories which include basic metadata, macro-based content features, DDE patterns, OLE object features, and XML structural features. The system aims to detect multiple malicious document characteristics through static analysis while maintaining lightweight functionality. The key components of the feature extraction phase are as follows:
  • Entropy & Size: Measures the level of randomness and the overall document size, often indicating embedded payloads or obfuscation.
  • Macros & VBA Keywords: These two components detect the presence of macros and extracts suspicious script keywords such as AutoOpen, Shell, CreateObject, and PowerShell.
  • DDE Detection: this component identifies the present of DDE patterns, which are known to execute external commands when the document is opened.
  • XML Structure Paths & Attributes: These components consider for Extracting the number of structural patterns and attributes in the internal XML trees (DOCX files are ZIP archives of XML data).
  • OLE Object Indicators: These features Count the embedded OLE objects and their types.
All extracted features are listed in Table III.
Table III: Summary of extracted features for Word document analysis.
Feature Group Purpose (Brief Description) No.
Basic Metadata Size, entropy (obfuscation signal) 2
Macro / VBA Indicators Macro presence, suspicious VBA keywords 2
DDE Detection DDE patterns that can trigger external execution 1
OLE Object Indicators Embedded OLE objects (counts/types) 2
XML Structure Paths & Attributes WordprocessingML structural attributes and path frequencies 36
Total 43
  • Note: No= Number of extracted Features
In the feature extraction phase, 43 static features were extracted from Word documents; however, not all of them were useful for classification task. Therefore, we applied the two-step feature selection process described earlier, using SHAP and Random Forest feature importance analysis to identify the most informative and lightweight features. The final set of 10 features was selected based on their high importance rankings in both SHAP and Random Forest evaluation. These features combine strong predictive capabilities with efficient computation which makes them appropriate for real-time malicious document detection systems. The final set of selected features for Word documents is listed in Table IV.
Table IV: Top 10 Selected Features for Word Document Using SHAP & Feature Importance.
No. Feature Name Description
1 ole_object_count Number of embedded OLE objects in the document
2 ole_object_type_count Number of distinct types of embedded OLE objects
3 macro_present Binary flag indicating the presence of embedded VBA macros
4 dde_present Detects presence of DDE command patterns that can trigger external process execution
5 vba_keywords_count Number of suspicious VBA macro keywords (e.g., createobject, powershell)
6 entropy Shannon entropy of the document’s raw content, indicating randomness or obfuscation
7 struct_ContentType Frequency of ContentType attributes in the DOCX internal XML
8 struct_PartName Count of structural PartName references used in the document package
9 file_size Size of the Word document in bytes
10 struct_pos Frequency of pos attributes in XML structure (layout positioning)
IV-B Excel
In order to analyze Excel attachment files statically, we created a complete feature extraction pipeline. This system identifies structural, content-based, and macro-related attributes associated with obfuscation and suspicious behavior in Excel documents. The extraction framework extracts 48 unique features from each file across five high-level categories, including file-level metadata, sheet/cell structure, cell content statistics, macro code metrics, and embedded resource/external interaction features.
In addition, the system uses Optical Character Recognition (OCR) to extract textual information from preview images embedded in Excel files which helps reveal hidden deceptive content. We excluded preview image text that embeds crawl hints (e.g., site watermarks).
The main categories of the extracted features are summarized as follows:
  • Entropy & Size: This group of features measures overall scale and text randomness that may indicate embedded payloads or obfuscation.
  • Cell Content & Encoding: These features identify what is inside cells and flags encoded data.
  • Macro Behavior & Code Metrics: This category captures VBA presence and obfuscation complexity in Excel files.
  • External References & System Interaction: This set of features indicates potential data exfiltration or code loading by using external links/APIs.
  • Structure & OCR Signals: The last category reveals any hidden components via layout and image-based lures.
Table V: Summary of Extracted Features for Excel Document Analysis.
Feature Group Purpose (Brief Description) No.
Basic Metadata Captures general file and worksheet dimensions 5
Cell Content Describes the content characteristics of spreadsheet cells 6
Formula and Hyperlinks Indicates spreadsheet complexity and external linking 2
Macro Structure Describes macro logic structure and complexity 11
Advanced Macro Features Captures deeper syntax patterns 7
Behavioral Indicators Detects interaction with system-level resources or external references 5
Sheet Properties Describes worksheet configurations 6
Preview Image Leverages embedded images for content-based cues 5
Total 48
  • Note: No= Number of extracted Features
All 48 extracted features from Excel files are summarized in Table V.
Although the initial set of 48 features was carefully extracted based on multiple factors, not all of them contributed equally to the classification model. Similar to the feature selection techniques applied for the previous file formats, we selected the top 10 most effective features that play a critical role in the detection phase. These selected features from Excel files are summarized in Table VI.
Table VI: Top 10 Selected Features for Excel Document Using SHAP & Feature Importance.
No. Feature Name Description
1 entropy_of_text Shannon entropy of extracted worksheet text; higher values suggest randomness/obfuscation or encoded payloads.
2 macro_chr_count Number of Chr/ChrW-style function uses in VBA, often used to construct strings and evade signatures.
3 macro_vocab_size Count of unique tokens in VBA macros; a proxy for macro code diversity/complexity.
4 macro_arithmetic_operator_count Number of arithmetic operators (e.g., +, -, *, /, Mod) used in macros; indicative of obfuscation logic.
5 macro_token_count Total token count in VBA macros; measures overall macro size/complexity.
6 macro_max_line_length Length of the longest line in VBA code; unusually long lines can hide concatenated or obfuscated payloads.
7 remote_template_present Binary flag indicating a reference to a remote template (external URL), which can enable external code loading.
8 numeric_cell_count Number of cells with numeric values across worksheets; helps profile content composition.
9 string_cell_count Number of cells with string values; large counts can correlate with embedded text or lures.
10 avg_cell_length Average character length of string cell contents; higher values may indicate embedded/encoded data.
IV-C PDF
To analyze PDF attachments statically, we designed a feature extraction pipeline focused on capturing low-level structural, behavioral, and metadata attributes. The system aimed to detect concealed malicious indicators which include hidden JavaScript code, exploitative actions, and obfuscated content that malicious PDFs typically use.
The framework extracts 40 features from each PDF file. These features are categorized into six main groups which include metadata & object statistics, structural & forms, stream & entropy analysis, and scripted actions & external references (JavaScript and URI). The system uses PyMuPDF, PyPDF2, and pdfminer, with an OCR-based fallback, in order to analyze different PDF types, including encrypted and image-only PDFs. The extracted feature categories are described as follows:
  • Stream Analysis and Entropy: These components calculate stream entropy and average stream size that appear between stream and end stream markers. The high entropy levels or large stream sizes show possible payload obfuscation.
  • JavaScript and Actions: The tool measures the number of occurrences for JavaScript, Launch, OpenAction, SubmitForm and URI entries (per-type counts) to identify auto-execution vectors which activate when files open or users interact with them.
  • External references & networking: The feature identifies any URL or action that uses non-standard ports which do not match 80 or 443. The function returns 1 when such a port exists but returns 0 when it does not.
  • Metadata & Object Statistics: The tool calculates file size information along with encryption flag status, object count, font object count and page number data as these metrics often correlate with suspicious manipulation or content embedding.
  • Obfuscation Indicators: These features evaluate name_obfuscations through encoded object names and searches for Filter[(nested filters) which might conceal payloads.
  • Fallback Text Recovery: It employs pdfminer, pdftotext and OCR as final extraction methods in order to achieve the highest possible document interpretability. The recorded text length serves as an indicator for both visibility and content density.
  • Structural & Forms: These features captured embedded content and interactive form structures that is able to be delivered with payload or user-driven execution paths.
Table VII: Summary of Extracted Features for PDF File Analysis.
Feature Group Purpose (Brief Description) No.
Basic Metadata Size, number of pages, encryption status, and metadata footprint 4
Content Properties Readable text and title length, embedded images, and OCR fallback signal 4
Stream Analysis Stream usage statistics, size, entropy, and number of object streams 5
Object Statistics Counts of objects, fonts, and cross-reference tables or entries 4
Embedded Files Presence and typical size of embedded attachments 2
Obfuscation Indicators Encoded names and nested filters that may conceal payloads 2
JavaScript / URI Indicators Script references and presence of external link indicators 3
Action Triggers Auto or interactive actions (e.g., Launch, OpenAction, AA) 5
Form Features Interactive form structures such as AcroForm or XFA 2
Encoding Techniques & Media Advanced codecs, compression, or media filters (e.g., JBIG2, RichMedia) 4
Behavioral Correlation Co-occurrence of multiple risky behaviors within the same object 1
Structure Validity Header validity and key trailer/startxref marker consistency 3
Non-Standard Usage Presence of URLs or actions using non-standard network ports 1
Total 40
  • Note: No = Number of Extracted Features
Similar to the feature extraction procedures applied to the previous file formats, 40 features were extracted from PDF files, followed by the selection of the top ten features that demonstrated the highest predictive impact during evaluation. These selected features emphasize metadata footprint, structural richness, stream activity/entropy, and header validity-signals that proved most discriminative in our evaluations. The complete set of extracted features is summarized in Table VII, while the selected top features for PDF files are listed in Table VIII.
Table VIII: Top 10 Selected Features for PDF Document Using SHAP & Feature Importance.
No. Feature Name Description
1 text_length Total number of text characters extracted; very low values can indicate image-only or obfuscated text.
2 total_filters Count of filter operators referenced in streams (e.g., /FlateDecode, /LZWDecode); proxies encoding complexity.
3 title_chars Number of characters in the document’s /Title metadata field.
4 file_size Size of the PDF file in bytes.
5 object_count Total number of indirect objects defined in the document.
6 stream_count Number of stream objects (binary data sections).
7 endstream_count Count of endstream markers closing stream objects (consistency with stream_count).
8 metadata_size Size in bytes of the /Metadata (XMP) stream; 0 if absent.
9 valid_pdf_header Binary flag indicating a valid %PDF- header with version token.
10 entropy_of_streams Average Shannon entropy of stream contents; higher values suggest compression/encryption/obfuscation.
IV-D HTML
The extraction of efficient attributes from HTML files requires a complete static analysis pipeline that considers both structural and content-level features. The proposed extraction system analyzes the entire HTML body structure which includes scripts and tags together with attributes and embedded links; therefore, it captures behavioral and obfuscation patterns which are typically used in phishing and other malicious HTML attachments by attacker.
By using this pipeline, 40 features were extracted from each HTML file that cover different behavioral categories such as content entropy, embedded JavaScript, tag composition, URL structure, and obfuscation techniques. The system aims to identify suspicious behaviors through static analysis without relying on dynamic execution. The extraction pipeline consists of the following essential components:
  • Entropy & Size Indicators: This set of features focuses on measuring document entropy and script block randomness, which often indicate obfuscation or payload encoding.
  • HTML Tag Composition: This feature summarizes the total number of tags by considering unique tag types, nesting depth, and embedded components that can be an evidence of anomalies.
  • JavaScript Analysis: Tracks the frequency of scripts and also the size of embedded JavaScript code as well as the presence of suspicious functions such as eval() and redirect patterns like window.location.
  • URL & Link Behavior: The features in this category considers multiple indicators that include external/internal link counts, base64-encoded segments, hostname digit ratios, average subdomain counts, suspicious URL characters that may indicate redirection or obfuscation.
  • Obfuscation & Suspicious Indicators: This set of features includes hidden iframes, high base64 occurrence count, hex encoding rates and suspicious words such as login, password, and secure. These words are commonly used by attackers in malicious HTML files.
All extracted features are summarized in Table IX.
Table IX: Summary of Extracted Features for HTML File Analysis.
Feature Group Purpose (Brief Description) No.
Basic Metadata Size, lines, overall entropy, whitespace formatting levels 5
Tag Structure Tag count/variety, nesting depth, comments, noscript/object usage 6
JavaScript Indicators Script presence/size/complexity, risky functions (e.g., eval) 7
Form Features HTML forms enabling data capture 1
Iframe Analysis Presence of iframes, including hidden ones 2
Redirection Indicators Automatic redirection mechanisms 1
Obfuscation Techniques Base64 usage, hex encoding, escaped characters 3
Suspicious Content Keywords Phishing keyword frequency and keyword-to-text ratio 2
URL and Link Analysis Link structure, lengths, subdomains, digits/punctuation signals 11
Image Analysis Image tag usage 1
Event-Based Behavior Suspicious event handlers (e.g., onload, onclick) 1
Total 40
Similarly, 40 features were extracted from HTML files, followed by the selection of the top 13 features that demonstrated the highest predictive impact during evaluation. The selected top 13 features for HTML files are presented in Table X.
Table X: Top 13 Selected Features for HTML Files Using SHAP & Feature Importance.
No. Feature Name Description
1 url_punct_char_count Total punctuation characters across all extracted URLs (e.g., / - = ? & : . _). Higher counts can indicate obfuscation or heavy query strings.
2 tag_count Total number of HTML tags in the document (overall structural volume/complexity).
3 whitespace_ratio Proportion of whitespace to non-whitespace in visible text; unusual spacing can signal templating or padding/obfuscation.
4 entropy Shannon entropy of the HTML source; higher values can indicate packed/obfuscated content.
5 form_count Number of <form> elements; often higher in credential-harvesting pages.
6 embedded_js_count Number of inline/embedded JavaScript blocks (e.g., <script> without external src).
7 html_whitespace_ratio Whitespace proportion computed over the raw HTML (including markup); complements whitespace_ratio which targets visible text.
[…]
9 min_link_length Minimum character length among extracted URLs; extreme short/long values can be indicative of shorteners or crafted links.
10 external_link_count Count of hyperlinks pointing to external hosts (e.g., <a href> with a different hostname).
[…]
12 internal_link_count Count of hyperlinks pointing within the same host (internal navigation density).
13 url_digit_count Total digit characters appearing across all extracted URLs; often higher in crafted/parameterized URLs.
IV-E Classification Phase
The features extracted in the previous sections were evaluated separately to determine their effectiveness in distinguishing malicious samples from benign ones across different data types. To perform this evaluation, we used three well-established machine learning classifiers including Random Forest, XGBoost and Decision tree classifiers. The reason for selecting these three models is their interpretability, robustness to feature scaling, and strong performance on structured datasets.
To evaluate both the extracted and selected features, the machine learning models were first trained using the complete feature set to assess overall performance, and then retrained using the selected subset to obtain a representative feature set suitable for lightweight models, making them more appropriate for deployment in resource-constrained environments and email gateways.
Figure 5: Proposed Quishing detection methodology and feature engineering.
QR code classification requires a separate approach, as these data are image-based and therefore demand image processing techniques for both feature extraction and classification. To perform both tasks efficiently, convolutional neural network (CNN) techniques are particularly convenient. Accordingly, a basic CNN-based model was employed to automatically identify discriminative image features for QR code analysis through automated learning instead of requiring human-designed features. After applying the CNN model, we also examined the URLs embedded in the QR codes by using discriminative lightweight language models, including BERT-Tiny, DeBERTa-v3, ModernBERT, and DeepSeek-R1 (Distill-Qwen), to assess their detection performance. The evaluation results are presented in the next section. Figure 5 illustrates the proposed Quishing detection methodology, encompassing two complementary approaches: one based on QR code decoding and URL tokenization, and the other utilizing CNN-driven image analysis.
V Experimental Results and Evaluation
@@ V-A Evaluation Metrics
[…]
V-B Evaluation Results
@@
V-B1 Word Doc Evaluation
In the feature extraction step, we extracted 43 static features from Word Document files, as described in section 4. To obtain a compact set of efficient and discriminative features, we selected top 10 features from SHAP and Random Forest feature importance rankings. These selected features were evaluated using three classifiers, and the detailed results are presented in Table XI. Additionally, Figure 9 (a) provides a visual representation of the results. The selected set of 10 features achieved 100% Accuracy, Precision, Recall, and F1-score, demonstrating the strong effectiveness of the selected feature subset. The confusion matrices for all three classifiers are shown in Figure 6. These matrices confirm the near-perfect classification performance of benign vs. malicious Word Document files by considering the top 10 selected features. Random Forest and XGBoost achieved perfect classification with no misclassifications, while the Decision Tree confusion matrix shows there is only a negligible error rate. These results confirm that the top 10 selected features enable highly accurate and efficient detection of malicious Word document files.
@@ V-B1 Word Doc Evaluation
V-B2 Excel Evaluation
In order to evaluate the extracted and selected Excel features, we performed the set of experiments that were conducted on Word documents. As explained before, a top 10 features were selected out of 48 extracted features and then the three machine learning classifiers were used for evaluation phase. Figure 9 (b) illustrates the four evaluation metrics, while Table XI presents their detailed results. Additionally, Figure 7 shows the confusion matrices of all three classifiers as well, which demonstrates the effectiveness of the extracted features.
@@ V-B2 Excel Evaluation
V-B3 PDF Evaluation
For PDF documents, a total of 40 features were initially extracted, as discussed in Section 4. Subsequently, the top 10 features were selected based on their importance scores. Similar to the previous file formats, these top 10 features were evaluated using the three classifiers described earlier. Table XI presents the detailed evaluation metrics. The results show that the selected features are highly effective, enabling accurate and efficient classification of benign and malicious PDF files. The results of these three models are provided in Figure 9 (c), which shows that all four evaluation metrics, Accuracy, Precision, Recall, and F1-score, exceed 99%.
@@ V-B3 PDF Evaluation
V-B4 HTML Evaluation
From each benign and malicious HTML sample, a set of 40 features were initially extracted, and subsequently, 13 features were selected that provided the best Accuracy, Precision, Recall and F1-score, while providing low training and testing time. By applying three classification models, XGBoost and Random Forest classifiers achieved approximately 93% in Accuracy, Precision, and F1-score, with a Recall of 92% which show the strength of selected features. The Decision Tree classifier achieved around 89% across all metrics, which still indicates satisfactory classification performance. The detailed results of these models for the selected HTML features are presented in Table XI and visualized in Figure 9 (d). The confusion matrices for all three classifiers are also illustrated in Figure 10.
Table XI: Per-classifier performance by data type.
@@ V-B4 HTML Evaluation
V-B5 QR code Evaluation
In order to evaluate the QR codes samples that were already generated from the malicious and benign URLs, we firstly followed an image processing approach by applying a basic CNN model to the QR code images. The corresponding performance results are provided in Table XII. As shown in this table, the CNN model accuracy is moderate, despite being trained on a large dataset, which proves that the image-based approach has some limitations in detection phase. To further investigate these limitations, additional visual and structural similarity analyses were performed. The results of these analyses are presented in Figure 11.
@@ V-B5 QR code Evaluation
The next step that we followed involved analyzing the URL strings corresponding to the QR codes to evaluate the effectiveness of benign and malicious URL samples in the detection phase from a lexical perspective. In order to examine different models, several lightweight language models (LLMs), including BERT-Tiny, DeBERTa-v3, ModernBERT, and DeepSeek-R1, were applied to the URL dataset in this experiment. The results presented in Table XII proves that LLM-based methods (URL lexical analyses) achieve better performance compared to the CNN model, as reflected by their higher Precision, Recall, and F1-score values. Moreover, the training time of these models is significantly lower than image based approach, since they process character level representations and tokenized text rather than images.
@@
Table XII: Performance results for malicious QR code detection using image-based approaches and recent lightweight LLMs applied to URL-based analysis.
[…]